Circular No. 6896

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980425 C. Tinney, R. Stathakis, and R. Cannon, Anglo-Australian Observa-

tory; and T. Galama, University of Amsterdam, report spectroscopy of the possible supernova in the galaxy ESO 184-G82 associated with the γ-ray burst GRB 980425 (IAUC 6895), obtained with the 3.9-m Anglo-Australian Telescope (+ 2dF spectrograph; range 620–720 nm, resolution 0.15 nm) on May 2.74118 UT. There is weak (EW = −0.18 nm) Hα emission, and possible S II emission indicating a galaxy redshift of z = 0.0085 ± 0.0002.

M. Wieringa, Australia Telescope National Facility (ATNF); D. A. Frail, National Radio Astronomy Observatory; S. R. Kulkarni, California Institute of Technology (CIT); J. L. Higdon and R. Wark, ATNF; J. S. Bloom, CIT; and the BeppoSAX GRB Team report: “Radio observations at 6 and 3 cm were made toward two x-ray sources in the field-of-view of GRB 980425 (IAUC 6884; GCN 61) with the Australian Telescope Compact Array starting on Apr. 28.73, 29.79, and May 5.55 UT. No radio sources were detected within the 1′ error radius of the two NFI sources. There is a bright radio source at α = 19^{h}35^{m}03^{s}.31, δ = −52^{o}50′44′′.7 (equinox 2000.0; ± 0′′.1). On Apr. 28 the flux densities of the source were 9 and 13 mJy at 6 and 3 cm, respectively, and were similar on Apr. 29 (9.9 and 13 mJy). On May 5 the flux densities of the source had increased dramatically, to 39 and 48 mJy at 6 and 3 cm, respectively. The position of the variable radio source coincides with the optical astrometric position given by Galama et al. on IAUC 6895. At the distance to ESO 184-G82, as implied by its redshift (Tinney et al., above), the radio source is already three times more luminous than SN 1988Z, one of the most luminous radio supernovae discovered, and is still brightening. Given the small likelihood of finding such an unusual radio source in the field, we suggest that the radio source may be related to GRB 980425.”

V4334 SAGITTARII S. Eyres, B. Smalley, and A. Evans, Keele University; and T. Geballe,

Joint Astronomy Center, Hilo, report infrared spectra obtained on Mar. 18: “There is strong He I absorption at 1.083 μm that was not present in 1997 July (Eyres et al. MNRAS, in press), indicating a source of high-energy photons. This may mean that the ejected envelope is dispersing, revealing the hot core. C I absorption at 1.069 μm, which was present in 1997 July, is no longer visible, but C_{2} lines have strengthened. Further observations, particularly of the light curve and optical spectrum, are urged.”

1998 May 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green